OpenStudy (anonymous):

Calculate the concentration of sodium ions made by diluting 45.0 mL of a 0.748 M solution of sodium sulfide (Na2S) to a total volume of 250.0 mL

OpenStudy (ookawaiioo):

Use M1V1 = M2V2 M = molarity V = volume

OpenStudy (anonymous):

I can't see, to get the correct answer using that formula. The right answer is 0.2693 M and I keep getting 1.346

OpenStudy (anonymous):

@Somy Can you help me figure out what I'm doing wrong?

OpenStudy (somy):

Step 1 M1V1 = M2V2 M = molarity V= make sure your unit is in dm3

OpenStudy (anonymous):

I'm not sure what dm3 means

OpenStudy (somy):

ml / 1000 = dm3 its a unit

OpenStudy (somy):

im trying to work it out

OpenStudy (anonymous):

That didn't come out right either. Hmmmm

OpenStudy (somy):

i know there is more to it im thinking

OpenStudy (anonymous):

If you can't figure it out I can always email my teacher. Thanks for trying :)

OpenStudy (somy):

well look after that you will get the molarity of Na2S but u need molarity of Na alone so you have the final molarity as 0.13464 M so this is total molarity now you have final molarity and volume so use it & find final mole of the solution so Molarity = mole/ dm3 mole = molarity x dm3 so molarity is 0.13464 M dm3 is 250/1000 and now u get whole mole

OpenStudy (anonymous):

That doesn't come up to .2693 M either

OpenStudy (somy):

after that you need mole of Na only which is why you do a some cross math stuff Na2S = molecular mass is (23*2) + 32 = get it now molecular mass of Na in there will be 23*2 = get it, because there are 2 of Na in this compound so mole of Na2S - molecular mass of Na2S mole of Na - molecular mass of Na atoms which is 46 now mole of Na = (mole of Na2S x 46) / molecular mass of Na2S and you got mole of Na now concentration of Na will be molarity = mole/ dm3 btw molarity & concentration are same thing so you have mole of Na and you have final dm3 = 250/1000 now use both values and get molarity of Na+ ions

OpenStudy (somy):

yeah, though the working is like this i cant seem to get the answer you say is the correct one

OpenStudy (anonymous):

Okay. Glad it's not just me. I'll email my teacher. Thank you!

OpenStudy (somy):

no problem :)

OpenStudy (somy):

nah, i checked it using your correct answer 0.2693M i did backward working and it brings me to a ridiculous number so i did this : molarity = mole/dm3 mole = 0.2693 x (250/1000) = 0.067325 mole of Na now i want to find mole of Na2S 0.067325 = 23 X = 78 X= 0.067325 x 78 /23 = 0.228 mole of Na2S so from here i'll get total conc of Na2S molarity = mole / dm3 molarity = 0.228 / (250/1000) = 0.912M = M2 woah thats like triple of the answer using M1V1 = M2V2 0.748 x (45/1000) = M2 x (250/1000) M2 = 0.13464 its a mistake, in your correct answer, i dont know where you got it from but its wrong how can u dilute something and get higher concentration instead? look initial concentration of solution is 0.748M and final you see it here as 0.912M it increased that cant be because the q states 'diluted' final answer of dilution of solution HAS to be less in concentration lol i guess now you can prove that the 'correct' answer is wrong backward working helps to figure it out :D

OpenStudy (somy):

i mean more than triple lol

OpenStudy (anonymous):

So @Somy I figured out how to get to the correct answer. You work it out like this 0.748M x 45.0 mL = C2 x 250 mL 33.66/250= C2 x 250/ 250 .13464= C2 .13464 x2 = C2 C2=.2693 M You multiple the answer by 2 because there is 2 sodium ions is each mole of sodium sulfide

OpenStudy (somy):

yeah but the faulty part here is the concentration that you are getting up there as .13464M is the concentration of diluted Na2S but the q is asking for Na+ concentration IN that diluted solution so you can't really multiply that by 2 well ask your teacher, i curious too of whats wrong here

OpenStudy (somy):

im*

OpenStudy (anonymous):

I did ask her. That's what she said needed to be done to it